OpenStudy (anonymous):

exact value of sin 5pi/12

jimthompson5910 (jim_thompson5910):

Hint: use the formula \[\large \sin\left(\frac{\theta}{2}\right) = \sqrt{\frac{1 - \cos(\theta)}{2}}\] in this case, \[\Large \theta = \frac{5\pi}{6}\]

OpenStudy (anonymous):

well i changed it so that it into degrees so that sin(45+35). i don't know where to go from here.

jimthompson5910 (jim_thompson5910):

45+35 = 80, which isn't the same as 5pi/12 since 5pi/12 = 75 degrees

OpenStudy (anonymous):

i meant 30

jimthompson5910 (jim_thompson5910):

oh you could do it like that sin(45 + 30) = sin(45)cos(30) + cos(45)sin(30)

jimthompson5910 (jim_thompson5910):

see where to go from here?

OpenStudy (anonymous):

i think so? thank you so much. i'll let you know if i have any more trouble :)

jimthompson5910 (jim_thompson5910):

alright great

OpenStudy (anonymous):

so i have gotten to